# Towards a Mechanistic Explanation of Diffusion Model Generalization

### v2

@@ 1 Introduction
Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021) have become the de facto standard for modelling image (Rombach et al., 2022) and video (Harvey et al., 2022) data due to their high sample quality and generalization abilities. When properly tuned, diffusion models produce samples which are distributionally similar to their training set, but are not exact copies of training data (Zhang et al., 2023).
This behaviour is remarkable, as linear increases in data dimension require exponentially more training samples to model the data density (Bellman, 1966). Avoiding this curse of dimensionality requires inductive biases which enable generalization from sparse examples (Goyal & Bengio, 2022). Recently, research has found that diffusion models produce near-identical samples despite differences in architecture, optimization, or diffusion hyperparameters (Zhang et al., 2023), suggesting there may be inductive biases common to all image diffusion models.
Figure 1: Denoiser outputs given shared reverse process noisy inputs. Column 1: Optimal empirical denoiser, i.e. what a “perfect” neural network denoiser would output if appropriately parameterized and trained to achieve minimal loss on the diffusion denoising loss (Equation 5). Columns 2-4: Outputs from various denoising neural networks. For $t<3.3$ all networks deviate from the optimal denoiser in similar ways. Column 5: Our learning-free patch set posterior composite denoiser produces qualitatively similar outputs to the neural network denoisers, suggesting that neural networks may generalize in part via patch denoising and composition.
Diffusion models generate samples through an iterative process in which noise is progressively removed through repeated denoising operations. Crucially, at each step of this process, there exists an optimal denoiser function which can be expressed as a simple weighted average of the training dataset (Vincent, 2011; Karras et al., 2022). However, using this optimal function in the denoising sampling procedure results in exact replication of the training dataset without generalization (Gu et al., 2023). The remarkable generalization capabilities of diffusion models therefore emerge from repeated neural network approximation errors relative to this optimal denoiser. These deviations from optimality, accumulated over the sampling procedure, compound to produce diverse samples.
@@
From this conclusion, we hypothesize that neural diffusion model generalization arises in part through locally biased operations. We establish evidence for this hypothesis by approximating these operations with patch-based empirical denoisers. Using these patch estimators, we demonstrate that for large portions of the forward diffusion process, local denoisers are equivalent to regions of the optimal denoiser. Further, we find that over the portion of the sampling procedure in which network denoisers deviate from optimal denoisers, patches of network outputs are closely approximated by patch empirical denoisers.
Finally, we propose our Patch Set Posterior Composites (PSPC) denoiser which aggregates patch empirical denoisers across varying spatial locations to approximate the hypothesized local mechanism of denoiser generalization. Comparing our fully training-free, empirical denoiser to network denoisers, we find PSPC and network denoisers are more similar to each other than to the optimal denoiser. Furthermore, samples produced using our denoiser share structural similarities to those produced by neural-network parameterized diffusion models. These findings provide strong evidence to conclude that patch denoising and composition comprise a significant portion of the generalization behaviour of image diffusion models.
@@ 2 Background
Diffusion models are based on a forward diffusion process which gradually add Gaussian noise to a data distribution $p(\mathbf{x}),\mathbf{x}\in\mathbb{R}^{d}$. This forward diffusion process can be described through stochastic differential equations of the form
@@
Hereafter, we will refer to the denoiser of Equation 6 as either the optimal denoiser.
@@ 3 Inductive Biases of Network Denoisers
Although Equation 6 is the optimal solution to the diffusion score matching objective, sampling using this denoiser can only reproduce exact copies of training data (Gu et al., 2023). The presence of generalization in image diffusion models beyond their training set therefore implies that denoiser networks make approximation errors to the optimal empirical denoiser. Further, the similarity of diffusion samples across several confounding factors (Zhang et al., 2023) suggests a shared class of approximation bias. To understand the generalization of diffusion models, we must understand and characterize these approximation errors.
To begin, we simply compare the denoiser outputs of four unconditional near-SoTA diffusion models trained on CIFAR-10 (Krizhevsky et al., 2009) to the optimal denoiser of that dataset. We evaluate models parameterized by NCSN++ (Song et al., 2021; Karras et al., 2022), DDPM++(Song et al., 2021; Karras et al., 2022), DiT (Peebles & Xie, 2023) and U-ViT (Bao et al., 2022) architectures.
Figure 2 plots the mean squared error (MSE) between network and optimal denoisers, evaluated over over 150 discrete $t\in[0.01,100]$, drawing 10,000 $\mathbf{z}$ samples from the forward process at each point. We observe that across all architectures, denoisers exhibit low error for both small and large values of $t$, but substantial error for $t\in[0.3,10]$, corroborating the findings of Niedoba et al. (2024). From the the right portion of Figure 2, only the middle $t=3$ row has significant differences between the optimal and network denoiser outputs [^1] The noise in U-ViT’s output for $t=30$ is due to the amplification problem for $\epsilon$-predictor networks (Karras et al., 2022). At this $t$, we observe that all three networks make qualitatively similar approximation errors, despite significant differences in architecture and training hyperparameters (see Appendix A). This observation expands upon those of Zhang et al. (2023), suggesting that the similarity of samples produced by generalizing diffusion models is a product of a similarity in the their denoiser outputs at every point in the diffusion process. Denoiser output similarities further suggest that denoiser approximation errors are not random optimization artifacts, but the result of a shared inductive bias common to all image network denoisers.
@@ 3.1 Network Denoiser Gradients
Figure 3: Left: Comparison of average square patch sizes required to capture 50, 75, and 95% of the total gradient sensitivity heatmap. Right: Gradient sensitivity heatmaps for DDPM++ denoiser on CIFAR-10 for output pixel (15,15) across varying $t$.
One potential inductive bias of network denoisers is local bias. We investigate potential local bias in network denoisers by measuring the sensitivity of the DDPM++ denoiser (Karras et al., 2019) to each input pixel through their gradients. For a specific output position $(x,y)$ and diffusion time $t$, we define the gradient sensitivity heatmap as
$\mathbf{G}(x,y,t)=\mathop{\mathbb{E}}_{\mathbf{z}\sim p_{t}(\mathbf{x}^{(i)},\mathbf{z})}\left[\sum_{c=1}^{3}\left\lvert\nabla_{\mathbf{z}_{c}}D_{\theta}(\mathbf{x},t)_{x,y,c}\right\lvert\right].$ (8)
$\mathbf{G}(x,y,t)$ captures the channel-averaged absolute gradient of the network denoiser output at pixel $(x,y)$ with respect to $\mathbf{z}$, a measure of the sensitivity of the network denoiser to each input pixel. In practice, we evaluate the expectation of Equation 8 using 10,000 $\mathbf{z}$ samples drawn from the forward process per $t$. We plot four such heatmaps in Figure 3
The right portion of Figure 3 confirms that network denoisers demonstrate strong local inductive bias. At $t=0.03$, the network denoiser is almost exclusively sensitive to the same spatial location as the output pixel. As $t$ increases, so does the area in which the input gradient is concentrated.
We analyze this local inductive bias quantitatively by measuring the average side length of a square patch centered at pixel $(x,y)$ required to capture a fixed percentage of $\sum_{i,j}\mathbf{G}(x,y,t)_{i,j}$. We plot this quantity in the left panel of Figure 3. For all $t$, the majority of the input gradient is centered within a square $15\times 15$ pixel region around the output pixel. Further, the strength of the local bias is anti-correlated with $t$. While a $13\times 13$ patch is required to capture 50% of $\sum_{i,j}\mathbf{G}(x,y,t)_{i,j}$ at $t=30$, the same percentage can be accounted for in a $3\times 3$ patch at $t=0.03$.
@@ 4 Local Denoising Mechanisms
Section 3 presents strong evidence that diffusion model denoisers deviate substantially and similarly from the optimal denoiser. Moreover, Figure 3 finds that network denoiser gradients show evidence of local inductive bias. However, Figure 2 also shows that network denoisers accurately estimate the empirical posterior mean for most $t$, despite the inherent global sensitivity of this quantity. To resolve this apparent contradiction, we hypothesize that network denoiser are primarily focused on local computations whose combined result is equivalent to the optimal denoiser for most values of $t$.
Figure 4: Left: Comparison of patch posterior means with varying patch sizes to corresponding patches of the optimal denoiser over forward process samples. For $t<1$, relatively small square patch posterior means exactly match the optimal denoiser. Right: Comparison of patch posterior means with varying patch sizes to DDPM++ denoiser patches on $\mathbf{z}$ drawn from the reverse process. Patch posterior means estimate the network denoiser patches better than the optimal denoiser for all $t<3$.
@@
It is particularly convenient to work with the set of square cropping matrices. We define $\mathbf{C}(x,y,s)$ as a square cropping matrix such that $\mathbf{C}(x,y,s)\mathbf{x}^{(i)}$ is the $s\times s$ pixel patch of $\mathbf{x}^{(i)}$ with upper left corner at pixel $(x,y)$. For a patch size $s$ and a square image of spatial size $h\times h$, we denote $\mathcal{C}_{s}=\left\{\mathbf{C}(x,y,s)\mid x\left\{0,\ldots,h-s\right\},y\in\left\{0,\ldots,h-s\right\}\right\}$ as the set of all such cropping matrices.
We have hypothesized that local bias in network denoisers are the result of local denoising operations which approximate the optimal denoiser. However, in general, $\mathbb{E}_{p_{\mathcal{D}}}[\mathbf{x}_{\mathbf{C}}\mid\mathbf{z}_{\mathbf{C}},t]\neq\mathbf{C}\mathbb{E}_{p_{\mathcal{D}}}[\mathbf{x}\mid\mathbf{z},t]$ because $p_{t}(\mathbf{x}^{(i)}_{\mathbf{C}}\mid\mathbf{z}_{\mathbf{C}})\neq p_{t}(\mathbf{x}^{(i)}\mid\mathbf{z}).$ Why then would we expect network denoisers to use local posterior mean estimates to estimate the global posterior mean?
@@
Figure 5: Our PSPC denoiser. First, $\mathbf{z}$ is decomposed into patches using a set of cropping matrices. For each patch, we compute the patch posterior mean via Equation 9. Resulting means are then combined into one image and normalized by the by the number of patches which which overlap each pixel. Although square patches are visualized, PSPC can be used with any set of cropping matrices.
If network denoisers utilize localized denoising mechanisms to approximate the optimal denoiser over the forward process, we would also expect this mechanism to be used for reverse process samples also. Although patch posterior means approximate optimal denoiser patches well over the forward process, from the right subplot of Figure 4 this is not true of $\mathbf{z}$ sampled from the reverse process. Instead, we find patch-based denoisers estimate network denoiser patches well for all $t<5$ and especially for $3\times 3$ patch denoisers when $t<0.3$. This provides further evidence that network denoisers may utilize patch-denoising operations to approximate the optimal denoiser.
@@ 5 Patch Set Posterior Composites
Figure 4 demonstrates that patch posterior means of varying sizes approximate patches of diffusion network denoisers well over forward and reverse processes. This finding motivates us to propose our methodology which combines patch posterior means at multiple spatial locations.
Figure 6: Comparison of denoisers against DDPM++ over forward and reverse processes. Across all $t$, PSPC denoiser are better estimators of the network denoiser than the optimal denoiser. In most cases PSPC-Flex provides the best estimate of network denoisers.
Formally, for an arbitrary set of cropping matrices $\mathcal{C}=\left\{\mathbf{C}_{1},\ldots,\mathbf{C}_{L}\right\}$ which denote these spatial locations, we define our Patch Set Posterior Composite (PSPC) method as
@@
The patch set posterior composite of Equation 11 estimates the patch posterior mean for every patch $\mathbf{C}\in\mathcal{C}$ of $\mathbf{z}$. The output is produced by summing each of these patch posterior means together before normalizing by $\left(\sum\mathbf{C}^{\top}\mathbf{C}\right)^{-1}$, the number of patches which overlap each pixel. This process is described in Figure 5. In practice, we use the fast nearest-neighbour score estimators of Niedoba et al. (2024) to estimate the patch posterior means of Equation 11.
Equation 11 is dependent on the choice of patch set $\mathcal{C}$. We introduce two possible choices of patch set below which represent two variants of our methodology.

### v3

@@ 1 Introduction
Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021) have become the de facto standard for modelling image (Rombach et al., 2022) and video (Harvey et al., 2022) data due to their high sample quality and generalization abilities. When properly tuned, diffusion models produce samples that are distributionally similar to their training set, but are not exact copies of training data (Zhang et al., 2023).
This behaviour is remarkable, as linear increases in data dimensionality require exponentially more training samples to model the data density (Bellman, 1966). Avoiding this curse of dimensionality requires inductive biases that enable generalization from sparse examples (Goyal & Bengio, 2022). Recently, research has found that diffusion models produce near-identical samples despite differences in architecture, optimization, or diffusion hyperparameters (Zhang et al., 2023), suggesting the presence of inductive biases common to all image diffusion models.
Figure 1: Denoiser outputs given shared reverse process noisy inputs from CIFAR-10. Column 1: Optimal empirical denoiser, i.e. what a “perfect” neural network denoiser would output if appropriately parameterized and trained to achieve minimal loss on the diffusion denoising loss (Equation 5). Columns 2-4: Outputs from various denoising neural networks. For $t<3.3$ all networks deviate from the optimal denoiser in similar ways. Column 5: Our learning-free patch set posterior composite denoiser produces qualitatively similar outputs to the neural network denoisers, suggesting that neural networks may generalize in part via patch denoising and composition.
Diffusion models generate samples through an iterative process in which noise is progressively removed via repeated denoising operations. Crucially, at each step of this process, there exists an optimal denoiser function that can be expressed as a simple weighted average of the training dataset (Vincent, 2011; Karras et al., 2022). However, using this optimal function in the denoising sampling procedure results in exact replication of the training dataset without generalization (Gu et al., 2023). The remarkable generalization capabilities of diffusion models therefore emerge from repeated neural network approximation errors relative to this optimal denoiser. These deviations from optimality, accumulated over the sampling procedure, compound to produce diverse samples.
@@
From this, we hypothesize that neural diffusion model generalization arises in part through locally biased operations. We establish evidence for this hypothesis by approximating these operations with patch-based empirical denoisers. Using these patch estimators, we demonstrate that for large portions of the forward diffusion process, local denoisers are equivalent to regions of the optimal denoiser. Further, we find that over the portion of the sampling procedure in which network denoisers deviate from optimal denoisers, patches of network outputs are closely approximated by patch empirical denoisers.
Finally, we propose our Patch Set Posterior Composites (PSPC) denoiser which aggregates patch empirical denoisers across varying spatial locations to approximate the hypothesized local mechanism of denoiser generalization. Comparing our fully training-free, empirical denoiser to network denoisers, we find PSPC and network denoisers are more similar to each other than to the optimal denoiser (Figure 1). Furthermore, samples produced using our denoiser share structural similarities to those produced by neural-network parameterized diffusion models. These findings provide strong evidence to conclude that patch denoising and composition comprise a significant portion of the generalization behaviour of image diffusion models. Our open-source implementation of PSPC and other denoisers is available at https://github.com/plai-group/pspc.
@@ 2 Background
Diffusion models are based on a forward diffusion process that gradually adds Gaussian noise to a data distribution $p(\mathbf{x}),\mathbf{x}\in\mathbb{R}^{d}$. This forward diffusion process can be described through stochastic differential equations of the form
@@
Hereafter, we refer to the denoiser of Equation 6 as the optimal denoiser.
@@ 3 Inductive Biases of Network Denoisers
Although Equation 6 is the optimal solution to the diffusion score matching objective, sampling using this denoiser can only reproduce exact copies of training data (Gu et al., 2023). The presence of generalization in image diffusion models beyond their training set therefore implies that denoiser networks make approximation errors relative to the optimal empirical denoiser. Further, the similarity of diffusion samples across several confounding factors (Zhang et al., 2023) suggests a shared class of approximation bias. To understand the generalization of diffusion models, we must understand and characterize these approximation errors.
To begin, we simply compare the denoiser outputs of four unconditional diffusion models trained on CIFAR-10 (Krizhevsky et al., 2009) to the optimal denoiser of that dataset. We evaluate models parameterized by NCSN++ (Song et al., 2021; Karras et al., 2022), DDPM++(Song et al., 2021; Karras et al., 2022), DiT (Peebles & Xie, 2023) and U-ViT (Bao et al., 2022) architectures [^1] For architecture details, see Appendix A..
Figure 2 plots the mean squared error (MSE) between network and optimal denoisers, evaluated over 150 discrete values of $t\in[0.01,100]$. For each $t$, we compute MSE over 10,000 $\mathbf{z}\sim p_{t}(\mathbf{z}|\mathbf{x})p_{\mathcal{D}}(\mathbf{x})$ drawn from the forward process. Across all architectures, we observe similar behaviour to Niedoba et al. (2024), Figure 3 - that network denoisers exhibit low MSE for both small and large values of $t$, but substantial error for $t\in[0.3,10]$. From the the right portion of Figure 2, only the middle $t=3$ row has significant differences between the optimal and network denoiser outputs [^2] The noise in U-ViT’s output for $t=30$ is due to the amplification problem for $\epsilon$-predictor networks (Karras et al., 2022).. At this $t$, we observe that all three networks make qualitatively similar approximation errors, despite significant differences in architecture and training hyperparameters (see Appendix A). This observation builds upon those of Zhang et al. (2023), suggesting that the similarity of samples produced by generalizing diffusion models is the product of corresponding similarities in denoiser outputs throughout the diffusion process. These denoiser output similarities further imply that denoiser approximation errors are not random optimization artifacts, but the result of a shared inductive bias common to all image network denoisers.
To focus our analysis and avoid redundancy, we hereafter restrict our attention to DDPM++ as a representative network denoiser. Additional quantitative and qualitative results for NCSN++, DiT, and U-ViT denoisers, confirming the consistency of our results across diverse architectures, are provided in Appendix D.
Figure 3: Left: Comparison of average square patch sizes required to capture 50%, 75%, and 95% of the total gradient sensitivity heatmap. As $t$ increases, larger patch sizes are required to capture a fixed percentage of the total gradient sensitivity heatmap. Right: Gradient sensitivity heatmaps for DDPM++ denoiser on CIFAR-10 for output pixel (15,15) across varying $t$.
@@ 3.1 Network Denoiser Gradients
One potential inductive bias (Goyal & Bengio, 2022) of network denoisers is local inductive bias, where denoiser outputs are more sensitive to spatially local perturbations of $\mathbf{z}$ than distant ones. We investigate potential local inductive bias in network denoisers by measuring the sensitivity of the DDPM++ denoiser (Karras et al., 2019) to each input pixel through their gradients. For each $t$, we define the gradient sensitivity heatmap as
$\mathbf{G}(x,y,t)=\mathop{\mathbb{E}}_{\mathbf{z}\sim p_{t}(\mathbf{x}^{(i)},\mathbf{z})}\left[\sum_{c=1}^{3}\left\lvert\nabla_{\mathbf{z}_{c}}D_{\theta}(\mathbf{z},t)_{x,y,c}\right\lvert\right].$ (8)
Here, $D_{\theta}(\mathbf{z},t)_{x,y,c}$ indicates the output of the network denoiser at spatial position $(x,y)$ and channel $c$ where where $x\in{1,\ldots,w},y\in\{1,\ldots,h\}$, and $c\in\{1,2,3\}$ for an RGB image of height $h$ and width $w$.
$\mathbf{G}(x,y,t)$ captures the channel-averaged absolute gradient of the network denoiser output at pixel $(x,y)$ with respect to $\mathbf{z}$, a measure of the sensitivity of an output pixel of the network denoiser to each input pixel. In practice, we evaluate the expectation of Equation 8 using 1,000 $\mathbf{z}$ samples drawn from the forward process per $t$. We plot four such heatmaps in Figure 3.
The right portion of Figure 3 confirms that network denoisers demonstrate strong local inductive bias. At $t=0.03$, the network denoiser is almost exclusively sensitive to the same spatial location as the output pixel. As $t$ increases, so does the area over which the input gradient is concentrated.
We quantify this local inductive bias by measuring the average side length of a square patch centered at pixel $(x,y)$ required to capture a fixed percentage of $\sum_{i,j}\mathbf{G}(x,y,t)_{i,j}$. We plot this quantity in the left panel of Figure 3. For all $t$, the majority of the input gradient is concentrated within a square $15\times 15$ pixel region around the output pixel. Further, the strength of the local inductive bias is inversely correlated with $t$. While a $13\times 13$ patch is required to capture 50% of $\sum_{i,j}\mathbf{G}(x,y,t)_{i,j}$ at $t=30$, the same percentage can be accounted for in a $3\times 3$ patch at $t=0.03$.
@@ 4 Local Denoising Mechanisms
Section 3 presents strong evidence that diffusion model denoisers deviate substantially and similarly from the optimal denoiser. Moreover, Section 3.1 finds that network denoiser gradients show evidence of local inductive bias. However, Figure 2 also shows that network denoisers accurately estimate the empirical posterior mean for most $t$, despite the inherent global sensitivity of this quantity. To resolve this apparent contradiction, we hypothesize that network denoisers perform local computations whose combined result is equivalent to the optimal denoiser for most values of $t$.
Figure 4: Left: Comparison of patch posterior means with varying patch sizes to corresponding patches of the optimal denoiser over forward process samples. For $t<1$, relatively small square patch posterior means exactly match the optimal denoiser. As $t$ increases, larger patch sizes are required to exactly estimate patches of the optimal denoiser. Right: Comparison of patch posterior means with varying patch sizes to DDPM++ denoiser patches on $\mathbf{z}$ drawn from the reverse process. Patch posterior means estimate the network denoiser patches better than the optimal denoiser for all $t<3$. As $t$ decreases, so does the patch size which best estimates network denoiser patches.
@@
Like the empirical denoiser of Equation 6, the patch posterior mean is a simple average over dataset elements, weighted by posterior probabilities. However, unlike the empirical denoiser which uses full images $\mathbf{x}^{(i)}$, the patch posterior mean is computed over a spatial subset of the dataset where each image is cropped by $\mathbf{C}$.
It is particularly convenient to work with the set of square cropping matrices. We define $\mathbf{C}(x,y,s)$ as a square cropping matrix such that $\mathbf{C}(x,y,s)\mathbf{x}^{(i)}$ is the $s\times s$ pixel patch of $\mathbf{x}^{(i)}$ with upper left corner at pixel $(x,y)$. For a patch size $s$ and a square image of spatial size $h\times h$, we denote $\mathcal{C}_{s}=\left\{\mathbf{C}(x,y,s)\mid x\left\{0,\ldots,h-s\right\},y\in\left\{0,\ldots,h-s\right\}\right\}$ as the set of all such cropping matrices. Notably, $\mathcal{C}_{s}$ does not include cropping matrices which would require any padding of $\mathbf{x}^{(i)}$ at the image boundaries.
We have hypothesized that local inductive biases in network denoisers are the result of local denoising operations which approximate the optimal denoiser. However, in general, $\mathbb{E}_{p_{\mathcal{D}}}[\mathbf{x}_{\mathbf{C}}\mid\mathbf{z}_{\mathbf{C}},t]\neq\mathbf{C}\mathbb{E}_{p_{\mathcal{D}}}[\mathbf{x}\mid\mathbf{z},t]$ because $p_{t}(\mathbf{x}^{(i)}_{\mathbf{C}}\mid\mathbf{z}_{\mathbf{C}})\neq p_{t}(\mathbf{x}^{(i)}\mid\mathbf{z}).$ Why then would we expect network denoisers to use local posterior mean estimates to estimate the global posterior mean?
@@
If network denoisers utilize localized denoising mechanisms to approximate the optimal denoiser over the forward process, we would expect this mechanism to be used for reverse process samples as well. Although patch posterior means approximate optimal denoiser patches well over the forward process, from the right subplot of Figure 4 this is not the case for $\mathbf{z}$ sampled from the reverse process. Instead, we find patch-based denoisers estimate network denoiser patches well for all $t<5$ and especially for $3\times 3$ patch denoisers when $t<0.3$. This provides further evidence that network denoisers may utilize local denoising operations such as patch posterior means to approximate the optimal denoiser.
@@ 5 Patch Set Posterior Composites
Figure 5: Our PSPC denoiser. First, $\mathbf{z}$ is decomposed into patches using a set of cropping matrices. For each patch, we compute the patch posterior mean via Equation 9. Resulting means are then combined into one image and normalized by the by the number of patches that overlap each pixel. Although square patches are visualized, PSPC can be used with any set of cropping matrices.
Figure 4 demonstrates that patch posterior means of varying sizes approximate patches of diffusion network denoisers well over both forward and reverse processes. This finding motivates our proposed methodology, which combines patch posterior means at multiple spatial locations.
Formally, given an arbitrary set of cropping matrices $\mathcal{C}=\left\{\mathbf{C}_{1},\ldots,\mathbf{C}_{L}\right\}$ which denote these spatial locations, we define our Patch Set Posterior Composite (PSPC) method as
@@
The patch set posterior composite of Equation 11 estimates the patch posterior mean for every patch $\mathbf{C}\in\mathcal{C}$ of $\mathbf{z}$. The output is produced by summing each of these patch posterior means together before normalizing by $\left(\sum\mathbf{C}^{\top}\mathbf{C}\right)^{-1}$, the number of patches which overlap each pixel. This process is described in Figure 5. In practice, we estimate patch posterior means using the fast nearest-neighbour score estimators of Niedoba et al. (2024).
The performance of Equation 11 is dependent on the choice of patch set $\mathcal{C}$. We introduce two possible choices of patch set below which represent two variants of our methodology.